OpenStudy (j2lie):

If the slope of the line that passes through the points (a,0) and (1,-2) is 1/2, what is the value of a? @jim_thompson5910

OpenStudy (freckles):

have you applied the slope formula?

OpenStudy (freckles):

\[m=\frac{y_2-y_1}{x_2-x_1} \]

OpenStudy (j2lie):

No. I do not know that to do that with this problem.

OpenStudy (freckles):

well you are given two points on a line and the slope ... which makes me think slope formula since only one of the values is unknown in the pairs given you are given m is 1/2 (x1,y1)=(1,-2) and (x2,y2)=(a,0)

OpenStudy (freckles):

\[\frac{1}{2}=\frac{0-(-2)}{a-1}\] solve for a

OpenStudy (j2lie):

a = -2

OpenStudy (freckles):

how did you get that?

OpenStudy (j2lie):

0 + 2 is 2 and a - 1 is -1

OpenStudy (freckles):

it might make it easier on you if you flip both sides \[\frac{2}{1}=\frac{a-1}{0-(-2)}\] see if you can solve now

OpenStudy (freckles):

by the way 2/1 is just 2

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ a}}\quad ,&{\color{blue}{ 0}})\quad % (c,d) &({\color{red}{ 1}}\quad ,&{\color{blue}{ -2}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\implies \cfrac{{\color{blue}{ -2}}-{\color{blue}{ 0}}}{{\color{red}{ 1}}-{\color{red}{ a}}}=\cfrac{1}{2}\)

OpenStudy (j2lie):

I am confused now.

OpenStudy (freckles):

confused on how to solve 2=(a-1)/2 for a?

OpenStudy (j2lie):

Yes.

OpenStudy (freckles):

are confused on the flipping part? confused on the equation we got?

OpenStudy (j2lie):

I am confused on solving with A.

OpenStudy (freckles):

solving with A?

OpenStudy (freckles):

what does that mean

OpenStudy (freckles):

you mean solve for a?

OpenStudy (j2lie):

The variable is A in the equation

OpenStudy (freckles):

don't we have a not A? \[\frac{1}{2}=\frac{0-(-2)}{a-1} \\ \\ 0-(-2) \text{ is the same as } 2 \\ \text{ you have } \\ \frac{1}{2}=\frac{2}{a-1}\]

OpenStudy (freckles):

you can flip both sides like I did earlier \[\frac{2}{1}=\frac{a-1}{2} \\ \text{ now } \frac{2}{1}=2 \\ 2=\frac{a-1}{2}\]

OpenStudy (freckles):

try to undo that division by 2 ...multiply 2 on both sides

OpenStudy (j2lie):

1=4/a-1?

OpenStudy (freckles):

well 2 times 2 is 4 not 1 and how did you put a-1 back in the denominator? also 2/2=1 not 4...

OpenStudy (freckles):

oh you went backwards

OpenStudy (j2lie):

it is 1/2 times 2

OpenStudy (freckles):

do you see this equation 2=(a-1)/2 I was asking you multiply this equation by 2 on both sides

OpenStudy (freckles):

this gives you 4=a-1

OpenStudy (j2lie):

how did you get 2?

OpenStudy (freckles):

@Jhannybean you can finish i have to leave I'm sorry

OpenStudy (j2lie):

I get it now.

OpenStudy (jhannybean):

\[2=\frac{a-1}{2} \implies 4 = a-1 \implies a= 4+1\]

OpenStudy (j2lie):

That is what I got.

OpenStudy (j2lie):

Thank you.